OpenStudy (anonymous):

Write in Interval Notation a) the open interval with endpoints 2 and 3 b) the closed interval with endpoints 2 and 3 c) the half-open interval with endpoints 2 and 3 that contains 2 but not 3 2) Represent the following sets of numbers in two ways: interval notation and as subset of the number line a) 1 < ( with underlined) x < (no underline) 15 b) -2 < (both arrows with underline) x < 2 c) -5 (no underline) < x (with underline) < -3

OpenStudy (anonymous):

I have already finished the questions on my own and I would like to make sure that they are correct.

OpenStudy (mathmate):

You're welcome to post your answers for a check. As a check-up, you would use [ or ] for closed, and ( or ) for open intervals. Some schools use ] and [ (i.e. pointing outwards) to represent open intervals. Also, \(\le\) means less than or equal to, while \(\ge\) means greater than or equal to. For example, the interval \(5\le x\lt 7\) would be written as [5,7).